OpenStudy (anonymous):

a bus is travelling with 52 passenger. when it arrives at a stop, "y" passengers get off & 4 get on. at the next stop, one-third of the passengers get off & 4 get on. there are now 25 passengers. find "y"

OpenStudy (anonymous):

after first stop the number of passengers is :52 -y +4 second stop: (52 -y +4) - (52 -y +4) * (1/3) + 4 = 25 so we have to solve : (52 -y +4) - (52 -y +4) * (1/3) + 4 = 25

OpenStudy (anonymous):

can you do it ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but why is (52-y+4) written twice?

OpenStudy (anonymous):

at the second stop there was (52 -y +4) passengers and you have been told that 1/3 of them get off so we have to subtract (52 -y +4) * (1/3)

OpenStudy (anonymous):

dont ignore the fact that it is multiplied by 1/3

OpenStudy (anonymous):

so what do we do?

OpenStudy (anonymous):

(52 -y +4) - (52 -y +4) * (1/3) + 4 = 25 solve this!

OpenStudy (anonymous):

(52 -y +4) - (52 -y +4) * (1/3)=25-4=21..then?

OpenStudy (anonymous):

...?